In a double slit interference experiment, the separation between the slits is 1.0 mm, the wavelength of light used is 5.0 × 10-7 m and the distance of the screen from the slits is 1.0 m. (a) Find the distance of the centre of the first minimum from the centre of the central maximum. (b) How many bright fringes are formed in one centimeter width on the screen?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Given that, d = 1 mm = 10^–3 m, D = 1 m. So, fringe with = Dλ/d = 0.5 mm. a) So, distance of centre of first minimum from centre of central maximum = 0.5/2 mm = 0.25 mm b) No. of fringes = 10 / 0.5 = 20.